OpenStudy (anonymous):

if f(x) = 1 over x^2+3 what is f(1)? what is f(p^2)

OpenStudy (zzr0ck3r):

\[f(x)=\frac{1}{x^2+3}\implies f(1)=\frac{1}{1^2+3}=?\]

OpenStudy (anonymous):

well i got the answet to that. but idk the other question

OpenStudy (anonymous):

the f(p^2)

OpenStudy (zzr0ck3r):

so all we did here was put a 1 where we saw a x right?

OpenStudy (anonymous):

yea

OpenStudy (zzr0ck3r):

now just put a p^2 where you see an x and note that \[(p^2)^2=p^4\]

OpenStudy (zzr0ck3r):

so we have \[f(p^2)=\frac{1}{(p^2)^2+3}=\frac{1}{p^4+3}\]

OpenStudy (zzr0ck3r):

understand?

OpenStudy (anonymous):

so as my answer i would put 1 over p^4+3.. or should i simplify it?

OpenStudy (zzr0ck3r):

you cant simplify it

OpenStudy (zzr0ck3r):

its just that

OpenStudy (anonymous):

on my homewrork it says to give the integer or simplified fraction

OpenStudy (anonymous):

oh ok

OpenStudy (zzr0ck3r):

they just want you to understand that \[f(anything)=\frac{1}{(anything)^2+3}\]

OpenStudy (anonymous):

thank you!!!! :D

OpenStudy (zzr0ck3r):

np